# ManiFoundation Model for General-Purpose Robotic Manipulation of Contact Synthesis with Arbitrary Objects and Robots

### v1

@@ I Introduction
Foundation models [1] that are trained on broad data are becoming versatile tools for numerous tasks, revolutionizing fields like natural language processing and computer vision. Large Language Models (LLMs) and large Vision-Language Models (VLMs) serve as vast repositories of knowledge that can potentially enhance robotic capabilities. For instance, LLMs have exhibited their superior capacity to decompose complex tasks such as meal preparation into a series of subtasks, leading to substantial enhancements in robot planning and reasoning processes. Recently, people have started to investigate LLMs or VLMs in robotics to develop robot foundation models. These works either leverage pretrained LLMs or VLMs as high-level planners to call a set of pretrained low-level skills [2, 3, 4], or directly train low-level robot actions on massive datasets [5, 6, 7]. While these models excel at high-level task planning and semantic understanding, they face substantial challenges when confronted with the extremely high complexity of 3D object geometry, deformations, force constraints, and the lack of comprehension of different robot manipulator topologies when deploying robots to actually perform low-level manipulations. In other words, the breadth and adaptability of low-level manipulation capabilities have received considerably less attention compared to the high-level task-planning abilities in robotics. It is thus essential to build a manipulation foundational model that equips robots with the requisite low-level manipulation proficiency to engage with the physical world and competently execute a diverse range of tasks.
@@
To build a manipulation foundation model, the core problem is to define a unified task formulation for all kinds of robot manipulation tasks, just as the auto-regressive prediction task in GPTs [8], the mask prediction task in BERT [9], and the vision-language contrastive learning task in CLIP [10]. This formulation must: 1) handle arbitrary objects, including articulated rigid objects, solid rigid objects, and deformable objects; 2) support a diverse range of manipulation tasks, like grasping, pulling, re-orientating, and folding; 3) adapt for cross-platform integration, accommodating diverse robot designs and configurations; 4) integrate seamlessly with LLMs/LVMs for effective task execution and long-horizon planning.
In this work, we propose a foundation model framework for low-level robotic manipulation, aiming to meet the above requirements. We formulate the manipulation foundation task as the contact synthesis task. Specifically, the foundation model $\mathcal{F}$ takes as input 1) object point cloud and physical attributes (such as friction coefficients, Young’s modulus, etc), 2) task descriptions represented as target motion, 3) robot manipulator descriptions as point clouds, 4) manipulation region masks if needed, for example, the task requires the robot to grasp the pan’s handle. The foundation model $\mathcal{F}$ outputs the contact points, contact forces or post-contact motion, so that robots can manipulate the object toward its target motion. We generate a large-scale annotated dataset to train the model. We test our ManiFoundation model over diverse experiments to verify its effectiveness both in simulation and the real world. In summary, our primary contributions include:
  • We propose a holistic framework towards foundation models for general-purpose robotic manipulation via contact synthesis, achieving broad generalization over various objects, robots, and tasks.
  • We implement the ManiFoundation model consisting of both neural network backbones for visual and physical feature extraction and an iterative contact wrench optimization method that generates feasible robot poses and avoids collisions and penetrations.
  • We develop a comprehensive pipeline to generate a large-scale dataset with annotations covering diverse objects and robot hands. This dataset will be made publicly available on our project website.
  • We conduct extensive experiments to verify the effectiveness of our proposed framework both in simulation and real-world experiments over various manipulation tasks including (articulated) rigid object manipulations and deformable object manipulations.
@@ II-A Foundation Models in Robotics
Foundation models in robotics [11] recently received a lot of attention. SayCan [2] grounds LLMs using value functions of pretrained skills. VoxPoser [3] produces affordances and constraints from LLMs and vision-language models to combine as 3D value maps for motion planners. Codes-as-policies [4] use LLMs to write robot policy code leveraging natural language instructions. RoboCat [12] is self-improving AI agent for robotics, learning tasks across various arms and autonomously generating new data to enhance its skills. Robotics Transformer (RT-1) [5] is trained over 130k real-world data containing over 700 tasks collected by 13 robots over 17 months. RT-2 [6] enhances vision-language-action models using web and robotic data. RT-X [7] gathers a dataset from 22 robots with 527 robotic skills and shows that scaling data improves the generalization capabilities between multiple robot platforms. Our work focus on developing the foundation model specifically designed for manipulation skills via contact synthesis to support general robotic manipulation with the broad generalization over objects including diverse deformable objects, robots, and tasks.
@@ II-B Learning Visual Affordances
Learning affordances [13, 14, 15] has received increasing attention for robotic manipulation, including grasping [16, 17, 18] and articulated object manipulation [19, 20, 21]. Where2Act [19] predicts the actionable point-on-point cloud for primitive actions such as pushing and pulling. Affordance Diffusion [22] synthesizes complex interactions (i.e., an articulated hand) with a given object. Vision-Robotic Bridge (VRB) [23] train a visual affordance model on large-scale videos with human behavior to predict contact points and a post-contact trajectory learned from human videos. Our model provides contact points for diverse robot hands and objects, including dexterous hands and deformable objects, together with contact forces and post-contact motions to guide robot low-level action execution.
@@ III ManiFoundation Model Design
A robotic manipulation task can be decomposed into two parts: a task-planning module that gives the object trajectory, and a manipulation module that guides robots to achieve this trajectory. For any given robot task, we assume LLMs/VLMs, other models, or differentiable simulators with digital twins, can convert it to a sequence of object point clouds as waypoint subgoals. Examples of this step for tasks in our experiments are provided in our website. Note this aspect is not the primary focus of us. Here we investigate to develop a universal model that guide robots to manipulate objects towards the target pose or shape. The model should generalize over arbitrary objects, robots, and tasks (target shape or motion).
We formulate our ManiFoundation model as a contact synthesis problem. Specifically, our foundation model $\mathcal{F}$ takes as input 1) manipulated object descriptions including object point cloud and physical properties, 2) task descriptions as target motion, 3) robot descriptions as point clouds, 4) manipulation region masks if needed, for example, the task requires the robot to grasp the pan’s handle. It outputs the contact points and contact forces so that robots can manipulate the object toward its target shape or motion. Formally, let a point cloud with $N$ points and normals on each point be $\mathcal{P}=\{(x_{1},n_{1}),\cdots,(x_{N},n_{N})\}\in\mathbb{R}^{N\times 6}$, where $x_{i}\in\mathbb{R}^{3}$ is the position and $n_{i}\in\mathbb{R}^{3}$ is the inward normal vector of the $i$-th point. Let $\mathcal{P}_{o}$ and $\mathcal{P}_{h}$ be the input object and robot hand point cloud respectively. Let =$\mathcal{M}=\{m_{i}\}_{i=1}^{N},m_{i}\in\mathbb{R}^{3}$ be the task motion that should be achieved on each point of the object. Let the physical feature of the object be $\Gamma$ and the region mask be $\mathcal{R}$. The manipulation foundation problem is defined as finding a contact solution $M_{\mathcal{C}},M_{\mathcal{S}}=\mathcal{F}(\mathcal{P}_{o},\mathcal{P}_{h},\mathcal{M},\mathcal{R},\Gamma)$ that satisfies:
$M_{\mathcal{C}},M_{\mathcal{S}}=\min\|\mathbf{f}(\mathcal{P}_{o},\mathcal{P}_{h},\Gamma,M_{\mathcal{C}},M_{\mathcal{S}})-\mathcal{M}\|,$ (1)
where $M_{\mathcal{C}}=\{c_{i}\}_{i=1}^{H},c_{i}\in\mathbb{R}^{3}$ is the predicted contact point positions with $H$ points on the object point cloud, and $M_{\mathcal{S}}=\{s_{i}\}_{i=1}^{H},s_{i}\in\mathbb{R}^{3}$ is the contact force corresponds to each contact point $c_{i}$. $\mathbf{f}$ stands for the forward dynamics function that outputs the object per-point motion given the contact points and forces. An overview of our foundation model is shown in Fig. 2. In the following parts, we describe the modules one by one.
@@ III-A The ManiFoundation Network
The whole network structure is illustrated in Figure 3. There are mainly two parts of our network: the Feature Extractor and the Conditional Variational Autoencoder (CVAE). The feature extractor is designed to encode information from the object and robot manipulator, while the CVAE module is designed to fuse extracted features and generate contact point and force heatmaps for execution. We introduce the details of each module in the following sections.
@@ III-A1 Feature Extractor
We encode the object point cloud and robot manipulator point cloud separately with a PointNet++ [24] $\phi_{o}$ for $\mathcal{P}_{o}$ and a PointNet [25] as $\phi_{h}$ for $\mathcal{P}_{h}$. However, to deal with the diversity in both objects and task requirements, we also design specific modules before these feature extractors.
@@
Our model accommodates a large variety of object types, each with distinct physical properties. These properties are crucial for object manipulation, as they indicate different manipulation strategies. For example, consider a handkerchief and a piece of card of the same shape. To spin the handkerchief, one needs to grasp the four corners. However, to rotate a card you can apply friction to other areas. This leads to totally different contact points and forces solutions, thus encoding the physical attributions is necessary to distinguish different kinds of objects.
@@
This design offers several benefits. Firstly, it can handle objects possessing a wide and variable array of physical properties. Secondly, if additional physical properties are needed to identify new object types, new MLPs can be seamlessly added to map these new properties to feature vectors, which are then combined with previous feature vectors and fed into the mean pooling layer.
@@
In practice, certain manipulation tasks have extra constraints that emphasize which part of an object can be manipulated. These constraints often stem from human preferences or environmental constraints. For instance, it’s generally preferred for robots to grasp the handle of a teapot or pan, and a robot cannot grasp the bottom of a cup when it is placed on the table.
To accommodate these constraints, we design a manipulation region mask to encode the region information, denoted as =$\mathcal{R}=\left\{{r_{i}}\right\}_{i=1}^{N}\in\mathbb{R}^{N}$. $r_{i}=1$ means the $i$-th point of $\mathcal{P}_{o}$ is allowed to contact, otherwise $r_{i}=0$. In practice, this mask can be assigned by humans or generated from environmental constraints. Our model is trained to predict contact points exclusively within the manipulation region mask.
For deformable objects, specifying the point-wise target motion can be challenging. Our model accommodates this by accepting the target point cloud of the deformable objects as input. A scene flow model [28] processes both the current and target point clouds to deduce the required target motion. In contrast, for articulated or rigid objects, determining the target motion from the current to the target point cloud is straightforward.
@@ III-A2 CVAE for Contact Point and Force/Motion
In practice, multiple contact point and force combinations can achieve the same target motion for an object. For instance, to advance a cube, one might push from behind or pull from its sides. To generate a range of solutions, we utilize a Conditional Variational Autoencoder (CVAE) for creating contact point heatmap and force/motion heatmap. The contact point heatmap, with dimensions $\mathbb{R}^{N\times 1}$, assigns a probability score to each point indicating its likelihood of being a contact point. The force/motion map, sized $\mathbb{R}^{N\times 3}$, represents either the contact force or the post-contact motion at each point.
@@ III-B Robotic Hand Pose Refinement
Given the contact points predicted by our network, we can directly manipulate deformable objects by simple grasping the contact points to achieve the target motion. However, for articulated and rigid object manipulation generally requires the coordinated effort of all fingers, while the contact points predicted by our network may not lead to feasible joint values of the robotic hand that satisfy the robot kinematics and avoid penetrations between robots and objects. Thus, we develop an optimization module for robotic hand pose refinement.
@@ III-B1 Pose Initialization
We infer the initial hand pose from the predicted contact points described in Sec. III-A2. Specifically, for a multi-fingered hand, we can iterate all possible matches between fingertips and the predicted contact points. For instance, with four contact points and a four-fingered robotic hand, there would be $4!=24$ possible matches. For each match, we consider the robot hand in its neutral joint poses as a rigid object and apply the Iterative Closest Point (ICP) algorithm to align robot fingertip points with contact points on the object. The result of ICP is the hand 6D pose for this match. Under this hand 6D pose, we generate a contact heatmap by applying a Gaussian kernel on the object point cloud at locations nearest to the fingertip positions. The best match, yielding the highest Intersection over Union (IoU) score between the generated heatmap and the predicted contact point heatmap, is selected. This initial 6D pose is then used for the robotic hand pose refinement.
@@ III-B2 Target Motion and Target Wrench
where object mass and inertia are denoted as $M$ and $I$, which are computed approximately by estimating the volume of the object. We normalize $\boldsymbol{\hat{w}}$ to target wrench $\boldsymbol{w}$ for practical simplicity. In practice, we care more about the direction of wrench and leave low-level force/torque control to take care of action to move objects towards target shapes.
@@ III-B3 Wrench Optimization
Given a target wrench and robot hand’s joint poses, we calculate the optimal contact forces $\boldsymbol{f}^{*}$ exerted by $m$ fingertips onto the object to minimize the norm of wrench residual $\boldsymbol{r}$ in Eqn. 7.
@@
$\boldsymbol{f}_{1:m}^{*}(\boldsymbol{w},\boldsymbol{q})=\underset{\boldsymbol{f}_{1:m}}{\arg\min}\;&\Big{\|}\boldsymbol{r}(\boldsymbol{w},\boldsymbol{q},\boldsymbol{f}_{1:m})\Big{\|}^{2},\\ \textrm{s.t. }&\boldsymbol{f}_{1:m}\in FC(\mu),\end{aligned}$ (7)
where $G_{1:m}$ is the grasp mapping matrices that maps 3d contact force at fingertips to 6d wrench applied to the object and $FC(\mu)$ is the friction cone with coefficient $\mu$ guaranteeing that contact forces satisfy Coulomb’s law [30]. We do not require fingertips to always touch the object’s surface through optimization. If fingertip $i$ is away from the surface, $G_{i}$ is calculated from its center position $\boldsymbol{p}_{i}$ and the normal vector at its projection onto the surface $\bar{\boldsymbol{n}}_{i}$.
The optimization problem that improves the hand pose $\boldsymbol{q}$ to enhance the grasp quality is formulated as:
@@
where $\boldsymbol{r}^{*}(\boldsymbol{w},\boldsymbol{q})=\boldsymbol{r}\big{(}\boldsymbol{w},\boldsymbol{q},\boldsymbol{f}_{1:m}^{*}(\boldsymbol{w},\boldsymbol{q})\big{)}$ is the minimum wrench residual. Our optimization objective in Eqn. 8 comprises two terms. The first term is the minimized norm of wrench residual, which measures the current grasp’s capability to induce the desired motion in the object. The second term is the aggregate distance between fingertips and the object’s surface, defined using the object’s signed distance field ($\Phi$) and the fingertip center positions $\boldsymbol{p}_{1:m}$ obtained from forward kinematics. for any 3D point $\boldsymbol{x}$, we locate the nearest points $\bar{\boldsymbol{p}}$ on the object’s surface and the outward normal $\bar{\boldsymbol{n}}$. $\Phi(\boldsymbol{p})=\bar{\boldsymbol{n}}\cdot(\boldsymbol{p}-\bar{\boldsymbol{p}})$, with $\frac{\partial\Phi(\boldsymbol{x})}{\partial\boldsymbol{x}}=\bar{\boldsymbol{n}}$. The second term in Eqn. 8 quantifies the sum of the distances between the fingertip centers and the object’s surface.
We also add a set of constraints to make the resulting hand pose practical. Eqn. 9 describes the robot joint limits and we want the system penetration-free. For simplicity, we use a set of $n$ spheres with different radii $\epsilon_{1:n}$ to represent the complex collision shape of robot hand. The spheres are attached to the fingers and palm, so the center positions $\boldsymbol{c}_{1:n}(\boldsymbol{q})$ are obtained from forward kinematics. Eqn. 10 guarantees no penetration between the hand and object and Eqn. 11 guarantees no self-penetration within the fingers and palm of the robot hand.
We solve this nonlinear optimization problem iteratively by performing Taylor expansion to Eqn. 8-11 in the neighborhood of $\boldsymbol{q}$, $\boldsymbol{f}_{1:m}^{*}$ and obtaining a convex optimization problem (second-order cone programming) which could be solved quickly by CVXPY [31] as follows:
@@
In the above equations, the Jacobians $\frac{\partial\boldsymbol{p}(\boldsymbol{q})}{\partial\boldsymbol{q}}$ and $\frac{\partial\boldsymbol{c}(\boldsymbol{q})}{\partial\boldsymbol{q}}$ are provided by the simulation Jade [32]. The exact formula of grasp mapping matrices and the corresponding gradients are derived in supplementary. Besides the constraints derived from the original problem Eqn. 8-11, we also add Eqn. 18 restricting $\delta\boldsymbol{q}$ and $\delta\boldsymbol{f_{1:m}}$ in the neighbourhood of zero, so that Taylor expansion is reasonable. We then update the joint values of the robot hand by $\boldsymbol{q}\leftarrow\boldsymbol{q}+\delta\boldsymbol{q}^{*}$. The convergence conditions are $\|\boldsymbol{r}^{*}(\boldsymbol{w},\boldsymbol{q})\|<10^{-6}$ and $|\Phi\big{(}\boldsymbol{p}_{1:m}(\boldsymbol{q})\big{)}|<0.005$. We set the maximum number of iterations as 100. The optimization is successful if it reaches convergence in 100 iterations, otherwise, it fails and terminates after 100 iterations. The successful output of these iterative convex optimization is the joint values of the robot hand and associated torques to generate wrench closest to target wrench.

### v2

@@ I Introduction
Foundation models [1] trained on broad data are becoming versatile tools for numerous tasks, revolutionizing fields like natural language processing and computer vision. Large Language Models (LLMs) and Vision-Language Models (VLMs) serve as vast repositories of knowledge that can enhance robotic capabilities. For instance, LLMs have demonstrated superior capacity in decomposing complex tasks, such as meal preparation, into subtasks, significantly improving robot planning and reasoning processes. Recently, researchers have started investigating LLMs and VLMs in robotics to develop robot foundation models. These efforts either leverage pretrained LLMs or VLMs as high-level planners to execute a set of pretrained low-level skills [2, 3, 4], or train low-level robot actions on massive datasets [5, 6, 7]. While these models excel at high-level task planning and semantic understanding, they face challenges with the complexity of 3D object geometry, deformations, force constraints, and comprehending different robot manipulator topologies during low-level manipulations. In other words, the breadth and adaptability of low-level manipulation capabilities have received less attention compared to high-level task planning in robotics. Thus, it is essential to build a manipulation foundational model that equips robots with the requisite low-level manipulation proficiency to engage with the physical world and competently execute a diverse range of tasks.
@@
To build a manipulation foundation model, the core problem is defining a unified task formulation for all robot manipulation tasks, similar to auto-regressive prediction in GPTs [8], mask prediction in BERT [9], and vision-language contrastive learning in CLIP [10]. This formulation must: 1) handle arbitrary objects, including articulated rigid objects, solid rigid objects, and deformable objects; 2) support a diverse range of manipulation tasks, such as grasping, pulling, re-orientating, and folding; 3) adapt to cross-platform integration, accommodating various robot designs and configurations; and 4) integrate seamlessly with LLMs/LVMs for effective task execution and long-horizon planning.
In this work, we propose a foundation model framework for low-level robotic manipulation to meet the above requirements. We formulate the manipulation foundation task as a contact synthesis task. Specifically, the foundation model $\mathcal{F}$ takes as input: 1) object point cloud and physical attributes (e.g., friction coefficients, Young’s modulus), 2) task descriptions as target motion, 3) robot manipulator descriptions as point clouds, and 4) manipulation region masks if required (e.g., grasping a pan’s handle). The foundation model $\mathcal{F}$ outputs the contact points, contact forces, or post-contact motion to enable the robot to manipulate the object towards its target motion. We generated a large-scale annotated dataset to train the model. We tested our ManiFoundation model in diverse experiments, verifying its effectiveness in both simulations and real-world applications. In summary, our primary contributions are:
  • We propose a comprehensive framework for foundation models in general-purpose robotic manipulation via contact synthesis, achieving broad generalization across various objects, robots, and tasks.
  • We implement the ManiFoundation model, which includes neural network backbones for visual and physical feature extraction and an iterative contact wrench optimization method to generate feasible robot poses while avoiding collisions and penetrations.
  • We develop a pipeline to generate a large-scale annotated dataset covering diverse objects and robot hands, which will be publicly available on our project website.
  • We conduct extensive experiments to verify the effectiveness of our framework in both simulations and real-world settings, covering various articulated rigid and deformable object manipulation tasks.
@@ II-A Foundation Models in Robotics
Foundation models in robotics [11] have recently garnered significant attention. SayCan [2] grounds LLMs using value functions of pretrained skills. VoxPoser [3] combines affordances and constraints from LLMs and vision-language models into 3D value maps for motion planners. Codes-as-policies [4] leverages LLMs to write robot policy code from natural language instructions. Robotics Transformer (RT-1) [5] was trained on 130k real-world data points, covering over 700 tasks collected by 13 robots over 17 months. RT-2[6] enhances vision-language-action models using web and robotic data. RT-X [7] collects a dataset from 22 robots with 527 robotic skills, showing that scaling data improves generalization across multiple robot platforms. Our work focuses on developing a foundation model specifically designed for manipulation skills via contact synthesis, supporting general robotic manipulation with broad generalization across diverse deformable objects, robots, and tasks.
@@ II-B Learning Visual Affordances
Learning affordances [12, 13, 14] has received increasing attention for robotic manipulation, including grasping [15, 16, 17] and articulated object manipulation [18, 19, 20]. Where2Act [18] predicts the actionable point-on-point cloud for primitive actions such as pushing and pulling. Vision-Robotic Bridge (VRB) [21] train a visual affordance model on large-scale videos with human behavior to predict contact points and a post-contact trajectory learned from human videos. Our model provides contact points for diverse robot hands and objects, including dexterous hands and deformable objects, together with contact forces and post-contact motions to guide robot low-level action execution.
@@ III ManiFoundation Model Design
A robotic manipulation task can be divided into two parts: a task-planning module that determines the object trajectory, and a manipulation module that guides the robot to achieve this trajectory. For any given task, we assume LLMs/VLMs, other models, or differentiable simulators with digital twins can convert it into a sequence of object point clouds as subgoals. Examples of this step for tasks in our experiments are provided on our website. Note that this aspect is not our primary focus. Here, we aim to develop a universal model that guides robots to manipulate objects towards the target pose or shape. The model should generalize across arbitrary objects, robots, and tasks (target shape or motion).
We present our ManiFoundation model as a solution to the contact synthesis problem. Specifically, the model $\mathcal{F}$ takes as input: 1) manipulated object descriptions, including the object point cloud and physical properties; 2) task descriptions as target motion; 3) robot descriptions as point clouds; and 4) manipulation region masks if needed (e.g., a pan’s handle). The model outputs contact points and contact forces to manipulate the object towards its target motion. Formally, let a point cloud with $N$ points and normals on each point be $\mathcal{P}={(x_{1},n_{1}),\cdots,(x_{N},n_{N})}\in\mathbb{R}^{N\times 6}$, where $x_{i}\in\mathbb{R}^{3}$ is the position and $n_{i}\in\mathbb{R}^{3}$ is the inward normal vector of the $i$-th point. Let $\mathcal{P}o$ and $\mathcal{P}h$ be the input object and robot hand point clouds, respectively. Let =$\mathcal{M}=\{m_{i}\}_{i=1}^{N},m_{i}\in\mathbb{R}^{3}$ be the task motion that should be achieved on each point of the object. Let the physical feature of the object be $\Gamma$ and the region mask be $\mathcal{R}$. The manipulation foundation problem is defined as finding a contact solution $M_{\mathcal{C}},M_{\mathcal{S}}=\mathcal{F}(\mathcal{P}_{o},\mathcal{P}_{h},\mathcal{M},\mathcal{R},\Gamma)$ that satisfies:
$M_{\mathcal{C}},M_{\mathcal{S}}=\min\|\mathbf{f}(\mathcal{P}_{o},\mathcal{P}_{h},\Gamma,M_{\mathcal{C}},M_{\mathcal{S}})-\mathcal{M}\|$ (1)
where $M_{\mathcal{C}}=\{c_{i}\}_{i=1}^{H},c_{i}\in\mathbb{R}^{3}$ is the predicted contact points with $H$ points on the object point cloud, and $M_{\mathcal{S}}=\{s_{i}\}_{i=1}^{H},s_{i}\in\mathbb{R}^{3}$ is the contact force corresponds to each contact point $c_{i}$. $\mathbf{f}$ stands for the forward dynamics function that outputs the object per-point motion given the contact points and forces. Figure 2 provides a pipeline overview. The following sections describe each module in detail.
@@ III-A The ManiFoundation Network
The whole network structure is illustrated in Figure 3. There are mainly two parts of our network: the Feature Extractor and the Conditional Variational Autoencoder (CVAE). The feature extractor is designed to encode information from the object and robot manipulator, while the CVAE module is designed to fuse extracted features to generate contact points and force heatmaps for execution. We introduce the details of each module in the following sections.
@@ III-A1 Feature Extractor
We encode the object point cloud and robot manipulator point cloud separately, using PointNet++[22] ($\phi_{o}$) for $\mathcal{P}_{o}$ and PointNet[23] ($\phi_{h}$) for $\mathcal{P}_{h}$. We chose PointNet for the manipulators due to its simpler architecture and fewer parameters, making it suitable for the limited variety of manipulators. In contrast, PointNet++ is better suited for extracting features from the more diverse and complex object point clouds. To address the diversity in both objects and task requirements, we also design specific modules before these feature extractors.
@@
Our model accommodates a wide variety of objects, each with distinct physical properties that are crucial for manipulation. These properties dictate different manipulation strategies. For example, a handkerchief and a card of the same shape require different approaches: spinning a handkerchief involves grasping the four corners, while rotating a card involves applying friction to other areas. These differences necessitate encoding physical attributes to distinguish between different types of objects, leading to varied contact points and force solutions.
@@
This design offers several benefits. First, it can handle objects with a wide range of physical properties. Second, if additional properties are needed to identify new object types, new MLPs can be easily added to map these properties to feature vectors. These new vectors are then combined with existing ones and fed into the mean pooling layer.
@@
Certain manipulation tasks have constraints that specify which parts of an object can be manipulated, often due to human preferences or environmental factors. For example, robots are typically preferred to grasp the handle of a teapot or pan and cannot grasp the bottom of a cup when it’s on a table. To address these constraints, we design a manipulation region mask, denoted as =$\mathcal{R}=\left\{{r_{i}}\right\}_{i=1}^{N}\in\mathbb{R}^{N}$. $r_{i}=1$. Here, $r_{i}=1$ indicates that the $i$-th point of $\mathcal{P}_{o}$ is allowed to contact, while $r_{i}=0$ means it is not. This mask can be assigned by humans or generated based on environmental constraints.
For deformable objects, specifying point-wise target motion is challenging. We use a scene flow model [26] to process both current and target point clouds to deduce the target motion. For articulated or rigid objects, determining the target motion is straightforward.
@@ III-A2 CVAE for Contact Point and Force/Motion
In practice, multiple contact points and force combinations can achieve the same target motion for an object. For instance, to advance a cube, one might push from behind or pull from its sides. To generate a range of solutions, we utilize a Conditional Variational Autoencoder (CVAE) for creating contact point heatmap and force/motion heatmap. The contact point heatmap, with dimensions $\mathbb{R}^{N\times 1}$, assigns a probability score to each point indicating its likelihood of being a contact point. The force/motion map, sized $\mathbb{R}^{N\times 3}$, represents either the contact force or the post-contact motion at each point. We chose CVAE instead of diffusion models for its simpler structure, faster inference speed, and multimodal generation capabilities as demonstrated in our experiments.
@@ III-B Robotic Hand Pose Refinement
Given the contact points predicted by our network, we can directly manipulate deformable objects by simply grasping the contact points to achieve the target motion. However, articulated and rigid object manipulation generally requires the coordinated effort of all fingers, while the contact points predicted by our network may not lead to feasible joint values of the robotic hand that satisfy the robot kinematics and avoid penetrations between robots and objects. Thus, we develop an optimization module for robotic hand pose refinement.
@@ III-B1 Pose Initialization
We infer the initial hand pose from the predicted contact points described in Sec.III-A. For a multi-fingered hand, we iterate all possible matches between fingertips and contact points (e.g., $4!=24$ matches for a four-fingered hand). For each match, we fix the robot hand in its neutral joint values as a rigid object and apply the Iterative Closest Point (ICP) algorithm to align the fingertips with the predicted contact points. The resulting hand 6D pose for each match is used to generate a contact heatmap, and the best match, based on the highest Intersection over Union (IoU) score between the generated and predicted heatmaps, is selected. This initial alignment provides a good starting pose for the robot palm, roughly aligning fingertip positions with predicted contact points, thereby shortening subsequent optimization time and improving success rates.
@@ III-B2 Target Motion and Target Wrench
where object mass and inertia are denoted as $M$ and $I$, which are computed approximately by estimating the volume of the object. We normalize $\boldsymbol{\hat{w}}$ to target wrench $\boldsymbol{w}$ for simplicity. In practice, we care more about the direction of the wrench and leave low-level force/torque control to take care of action to move objects towards target shapes.
@@ III-B3 Wrench Optimization
Given a target wrench and robot hand’s joint poses, we calculate the optimal contact forces $\boldsymbol{f}^{*}$ exerted by $m$ fingertips onto the object to minimize the norm of wrench residual $\boldsymbol{r}$ in Eqn. 6.
@@
$\boldsymbol{f}_{1:m}^{*}(\boldsymbol{w},\boldsymbol{q})$ $=\underset{\boldsymbol{f}_{1:m}}{\arg\min}\;\Big{\|}\boldsymbol{r}(\boldsymbol{w},\boldsymbol{q},\boldsymbol{f}_{1:m})\Big{\|}^{2},\;$ (7)
[…]
where $G_{1:m}$ is the grasp mapping matrices that maps 3d contact force at fingertips to 6d wrench applied to the object. $FC(\mu)$ is the friction cone with coefficient $\mu$ guaranteeing that contact forces satisfy Coulomb’s law [28]. During optimization, fingertips are not required to always be in contact with the object surface. If fingertip $i$ is away from the surface, $G_{i}$ is calculated from its position $\boldsymbol{p}_{i}$ and the normal vector at its projection onto the surface $\bar{\boldsymbol{n}}_{i}$.
The optimization problem to improve grasp quality is:
@@
where $\boldsymbol{r}^{*}(\boldsymbol{w},\boldsymbol{q})=\boldsymbol{r}\big{(}\boldsymbol{w},\boldsymbol{q},\boldsymbol{f}_{1:m}^{*}(\boldsymbol{w},\boldsymbol{q})\big{)}$ is the minimum wrench residual. Our optimization objective in Eqn. 9 comprises two terms. The first term is the minimized norm of wrench residual, which measures the current grasp’s capability to generate the desired object motion. The second term is the sum of the distance between fingertips and the object’s surface, calculated using the object’s signed distance field ($\Phi$) and the fingertips’ positions $\boldsymbol{p}_{1:m}$ obtained from forward kinematics. Specifically, for any 3D point $\boldsymbol{x}$, we locate the nearest points $\bar{\boldsymbol{p}}$ on the object’s surface and the outward normal $\bar{\boldsymbol{n}}$. $\Phi(\boldsymbol{p})=\bar{\boldsymbol{n}}\cdot(\boldsymbol{p}-\bar{\boldsymbol{p}})$, with $\frac{\partial\Phi(\boldsymbol{x})}{\partial\boldsymbol{x}}=\bar{\boldsymbol{n}}$.
We also add a set of constraints to make the resulting hand pose practical. Eqn. 10 describes the robot joint limits. For simplicity, we use a set of $n$ spheres with different radius $\epsilon_{1:n}$ to represent the collision shape of the robot hand. The spheres are attached to the fingers and palm, so the centers’ positions $\boldsymbol{c}_{1:n}(\boldsymbol{q})$ can be obtained from forward kinematics. Eqn. 11 guarantees no penetration between the hand and object and Eqn. 12 guarantees no self-penetration.
We solve this nonlinear optimization problem iteratively by performing Taylor expansion to Eqn. 9-12 in the neighborhood of $\boldsymbol{q}$, $\boldsymbol{f}_{1:m}^{*}$ and obtaining a convex optimization problem (second-order cone programming) which could be solved quickly by CVXPY [29] as follows:
@@
In the above equations, the Jacobians $\frac{\partial\boldsymbol{p}(\boldsymbol{q})}{\partial\boldsymbol{q}}$ and $\frac{\partial\boldsymbol{c}(\boldsymbol{q})}{\partial\boldsymbol{q}}$ are provided by the simulation Jade [30]. The exact formula of grasp mapping matrices and the corresponding gradients are derived in supplementary. Besides the constraints derived from the original problem Eqn. 9-12, we also add Eqn. 19 restricting $\delta\boldsymbol{q}$ and $\delta\boldsymbol{f_{1:m}}$ in the neighborhood of zero, so that Taylor expansion is reasonable. We then update the joint values of the robot hand by $\boldsymbol{q}\leftarrow\boldsymbol{q}+\delta\boldsymbol{q}^{*}$. The convergence conditions are $\|\boldsymbol{r}^{*}(\boldsymbol{w},\boldsymbol{q})\|<10^{-6}$ and $|\Phi\big{(}\boldsymbol{p}_{1:m}(\boldsymbol{q})\big{)}|<0.005$. We set the maximum number of iterations as 100. The optimization is successful if it reaches convergence within 100 iterations, otherwise, it fails and terminates after 100 iterations. The successful output of this iterative convex optimization is the joint values of the robot hand and associated torques to generate the wrench closest to the target wrench.